# An ordering for the strength of functional dependence

### v1

@@ B.6 Proofs of Section 4
(i): Due to Theorem 1.6, we know that $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ implies $(F_{Y}\circ q_{Y\mid\mathbb{X}}^{\uparrow U}(V),U)\leq_{c}(F_{Y^{\prime}}\circ q_{Y^{\prime}\mid\mathbb{X}^{\prime}}^{\uparrow U}(V),U)$ and thus $\mu(Y,\mathbb{X})=\mu(F_{Y}\circ q_{Y\mid\mathbb{X}}^{\uparrow U}(V),U)\leq\mu(F_{Y^{\prime}}\circ q_{Y^{\prime}\mid\mathbb{X}^{\prime}}^{\uparrow U}(V),U)=\mu(Y^{\prime},\mathbb{X}^{\prime})$.
(ii): $\mathbb{X}$ and $Y$ are independent if and only if $q_{Y|\mathbb{X}}^{\uparrow U}(V)$ and $U$ are independent; see Remark 2.8(b). Since $\mu$ characterizes independence on $\mathcal{R}^{\uparrow}$, the latter is equivalent to $\mu(Y,\mathbb{X})=\mu(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)=0$.
@@
(iii): $Y$ perfectly depends on $\mathbb{X}$ if and only if $q_{Y|\mathbb{X}}^{\uparrow U}(V)$ and $U$ are comonotone; see Remark 2.8(b). Since $\mu$ characterizes comonotonicity on $\mathcal{R}^{\uparrow}$, the latter is equivalent to $\mu(Y,\mathbb{X})=\mu(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)=1$.
(iv): By Theorem 1.4, we have $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y,(\mathbb{X},\mathbb{Z}))$. Hence, (i) implies the statement.
@@
(v): Assume that $\mu(Y,\mathbb{X})=\mu(Y,(\mathbb{X},\mathbb{Z}))$. By information monotonicity of $\preccurlyeq_{ccx}$ and due to Theorem 1.6, we know that $(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)\leq_{c}(F_{Y}\circ q_{Y|(\mathbb{X},\mathbb{Z})}^{\uparrow U}(V),U)$. This implies $\mu(Y,\mathbb{X})=\mu(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)\leq\mu(F_{Y}\circ q_{Y|(\mathbb{X},\mathbb{Z})}^{\uparrow U}(V),U)=\mu(Y,(\mathbb{X},\mathbb{Z}))=\mu(Y,\mathbb{X})$ and thus $\mu(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)=\mu(F_{Y}\circ q_{Y|(\mathbb{X},\mathbb{Z})}^{\uparrow U}(V),U)$. Since $\mu$ is strictly increasing on $\mathcal{R}^{\uparrow}$, it follows that $(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)=_{c}(F_{Y}\circ q_{Y|(\mathbb{X},\mathbb{Z})}^{\uparrow U}(V),U)$ and thus, by Theorem 1.6, $(Y,\mathbb{X})=_{ccx}(Y,(\mathbb{X},\mathbb{Z}))$. However, the latter means that $Y$ and $\mathbb{Z}$ are conditionally independent given $\mathbb{X}$. For the reverse direction, assume that $Y$ and $\mathbb{Z}$ are conditionally independent given $\mathbb{X}$. Then $(Y,\mathbb{X})=_{ccx}(Y,(\mathbb{X},\mathbb{Z}))$ and thus, by Theorem 1.6, $(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(F_{Y}\circ q_{Y|(\mathbb{X},\mathbb{Z})}^{\uparrow U}(V),U)$. This implies $\mu(Y,\mathbb{X})=\mu(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)=\mu(F_{Y}\circ q_{Y|(\mathbb{X},\mathbb{Z})}^{\uparrow U}(V),U)=\mu(Y,(\mathbb{X},\mathbb{Z}))$. $\blacksquare$
@@ B.7 Proofs of Section 5
$(q_{Y|X}^{\uparrow U}(V),U)\leq_{c}(q_{Y^{\prime}|X^{\prime}}^{\uparrow U}(V),U)$ (59)
@@
$q_{Y|X}^{\uparrow U}(V)=F_{U}^{-1}(V)=F_{Y|X=q_{X}(U)}^{-1}(V),$ (60)
with $q_{Y|X}^{\uparrow u}(v)$ defined by (10). Similarly, we obtain $q_{Y^{\prime}|X^{\prime}}^{\uparrow U}(V)=F_{Y^{\prime}|X^{\prime}=q_{X^{\prime}}(U)}^{-1}(V)$. Now, observe that
$(Y,X)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(F_{Y|X=q_{X}(U)}^{-1}(V),q_{X}(U))\quad\text{and}\quad(Y^{\prime},X^{\prime})\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(F_{Y^{\prime}|X^{\prime}=q_{X^{\prime}}(U)}^{-1}(V),q_{X^{\prime}}(U));$ (61)
see (38). Since the concordance order is invariant under increasing transformations, we obtain from $q_{X}=q_{X^{\prime}}$ with (59) and (60) that $(Y,X)\leq_{c}(Y^{\prime},X^{\prime})$.
To proof the reverse direction, assume that $(Y,X)\leq_{c}(Y^{\prime},X^{\prime})$. Using (61) and (60), it follows that
$(q_{Y|X}^{\uparrow U}(V),q_{X}(U))\leq_{c}(q_{Y^{\prime}|X^{\prime}}^{\uparrow U}(V),q_{X^{\prime}}(U)).$ (62)
Since $Y\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\prime}$ and $X\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X^{\prime}\,$, (62) is equivalent to
$P(q_{Y|X}^{\uparrow U}(V)\leq y,U\leq F_{X}(x))\leq P(q_{Y^{\prime}|X^{\prime}}^{\uparrow U}(V)\leq y,U\leq F_{X}(x))$ (63)
for all $x,y\in\mathbb{R}$; see Proposition A.1(i) and B.2. Since $q_{Y|X}^{\uparrow u}(v)$ and $q_{Y^{\prime}|X^{\prime}}^{\uparrow u}(v)$ depend on $u$ only through $q_{X}(u)$, it is not difficult to see that (63) is equivalent to $P(q_{Y|X}^{\uparrow U}(V)\leq y,U\leq u)\leq P(q_{Y^{\prime}|X^{\prime}}^{\uparrow U}(V)\leq y,U\leq u)$ for all $y\in\mathbb{R}$ and $u\in(0,1)$. Using $Y\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\prime}$, this implies $(F_{Y}\circ q_{Y|X}^{\uparrow U}(V),U)\leq_{c}(F_{Y^{\prime}}\circ q_{Y^{\prime}|X^{\prime}}^{\uparrow U}(V),U)$ which is by Theorem 1.6 equivalent to $(Y,X)\leq_{ccx}(Y^{\prime},X^{\prime})$. $\blacksquare$

### v2

@@ B.6 Proofs of Section 4
For the proof of Theorem 4.4, we make use of the supermodular order $\mathbb{U}\leq_{sm}\mathbb{V}$ which is defined for bounded, bivariate random vectors $\mathbb{U}=(U_{1},U_{2})$ and $\mathbb{V}=(V_{1},V_{2})$ by
$\mathbb{E}f(\mathbb{U})\leq\mathbb{E}f(\mathbb{V})\quad\text{for all supermodular functions }f.$ (B.27)
Recall that $f\colon\mathbb{R}^{2}\to\mathbb{R}$ is said to be supermodular (submodular) if $f(\mathbb{x})+f(\mathbb{y})\leq$ $(\geq)$ $f(\mathbb{x}\wedge\mathbb{y})+f(\mathbb{x}\vee\mathbb{y})$ for all $\mathbb{x},\mathbb{y}\in\mathbb{R}^{2}$. For bivariate random vectors, the supermodular order and the concordance order are equivalent. For bivariate random vectors from the same Fréchet class, the supermodular order is even equivalent with the pointwise order of distribution functions as follows; see [44, Theorem 2.5].
Lemma B.6 (Characterization of supermodular order)
If $\mathbb{U}=(U_{1},U_{2})$ and $\mathbb{V}=(V_{1},V_{2})$ have equal marginal distributions, then the following statements are equivalent.
  1. (i)
    $\mathbb{U}\leq_{sm}\mathbb{V}$,
  2. (ii)
    $P(U_{1}\leq u_{1},U_{2}\leq u_{2})\leq P(V_{1}\leq u_{1},V_{2}\leq u_{2})$ for all $(u_{1},u_{2})\in\mathbb{R}^{2}$.
Next, recall that, for submodular cost functions $c$, the comonotone coupling solves the optimal transport problem
$\mathcal{W}_{c}(\nu,\nu^{\prime}):=\inf_{\gamma\in\Pi(\nu,\nu^{\prime})}\int c(y,y^{\prime})\mathsf{\,d}\gamma(y,y^{\prime})$ (B.28)
as follows (see [49, Theorem 3.1.2]).
Lemma B.7 (Comonotone Coupling)
Let $\nu$ and $\nu^{\prime}$ be distributions on $\mathbb{R}$, and let $Z\sim\nu$ and $Z^{\prime}\sim\nu^{\prime}$ be random variables. Assume that the cost function $c$ is submodular. Then, for $V$ uniform on $(0,1)$, the comonotone coupling $(F_{Z}^{-1}(V),F_{Z^{\prime}}^{-1}(V))$ solves the optimal transport problem (B.28), i.e. $\mathcal{W}_{c}(\nu,\nu^{\prime})=\mathbb{E}c(F_{Z}^{-1}(V),F_{Z^{\prime}}^{-1}(V))$.
We also make use of the following quantile-based information monotonicity.
Lemma B.8 (A variant of information monotonicity)
For a random variable $W$ and random vector $(\mathbb{X},\mathbb{Z})$, let $V\sim U(0,1)$ be independent of $(\mathbb{X},\mathbb{Z})$. Then $F_{W|\mathbb{X}}^{-1}(V)-F_{W}^{-1}(V)\leq_{cx}F_{W|(\mathbb{X},\mathbb{Z})}^{-1}(V)-F_{W}^{-1}(V)$.
By information monotonicity of the conditional convex order, we have $(W,\mathbb{X})\preccurlyeq_{ccx}(W,(\mathbb{X},\mathbb{Z}))$, see Theorem 1.4. Then we obtain from Lemma B.4 for all $y,y^{\prime}\in\mathbb{R}$ that
$P(F_{W|\mathbb{X}}^{-1}(V)\leq y,$ $F_{W}^{-1}(V)\leq y^{\prime})=\mathbb{E}\min\{F_{W|\mathbb{X}}(y),F_{W}(y^{\prime})\}$
$\geq\mathbb{E}\min\{F_{W|(\mathbb{X},\mathbb{Z})}(y),F_{W}(y^{\prime})\}=P(F_{W|(\mathbb{X},\mathbb{Z})}^{-1}(V)\leq y,F_{W}^{-1}(V)\leq y^{\prime})\},$
where the inequality follows from the version of the conditional convex order in Remark 1.3a. Since $F_{W|\mathbb{X}}^{-1}(V)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}F_{W|(\mathbb{X},\mathbb{Z})}^{-1}(V)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}F_{W}^{-1}(V)$ by Lemma B.4, it follows from Lemma B.6 that
$(F_{W|\mathbb{X}}^{-1}(V),F_{W}^{-1}(V))\geq_{sm}(F_{W|(\mathbb{X},\mathbb{Z})}^{-1}(V),F_{W}^{-1}(V)).$ (B.29)
Now, the statement follows from [58, Theorem 9.A.18]. $\blacksquare$
First we show that the denominator in (31) is positive. Therefore, let $W$ and $W^{\prime}$ be independent random variables both with distribution $\nu$. Applying Jensen’s inequality twice, we obtain
$\int_{\mathbb{R}}\int_{\mathbb{R}}c(y,y^{\prime})\mathsf{\,d}\nu(y)\mathsf{\,d}\nu(y^{\prime})=\mathbb{E}h(W^{\prime}-W)\geq\int h(y^{\prime}-\mathbb{E}W)\mathsf{\,d}\nu(y^{\prime})>h(\mathbb{E}W^{\prime}-\mathbb{E}W)=0,$
where we use for the strict inequality on the one hand that $h$ is strictly convex at 0. On the other hand, we use that $Y$, and hence $W\stackrel{{\scriptstyle\mathrm{d}}}{{=}}F_{Y}(Y)$, is non-degenerate and that $W\stackrel{{\scriptstyle\mathrm{d}}}{{=}}W^{\prime}$. The last equality follows from the assumption $h(0)=0$.
For the rest of the proof, we will only consider the numerator of (31) because the denominators of $\overrightarrow{\mathcal{W}}_{c}(Y,\mathbb{X})$ and $\overrightarrow{\mathcal{W}}_{c}(Y^{\prime},\mathbb{X}^{\prime})$ coincide under the marginal constraint.
To prove (33), we need to show that the numerator of (31) is $\preccurlyeq_{ccx}$-increasing. We abbreviate $W:=F_{Y}(Y)$ and $W^{\prime}:=F_{Y^{\prime}}(Y^{\prime})$. Note that $W\stackrel{{\scriptstyle\mathrm{d}}}{{=}}W^{\prime}$ by the marginal constraint. Since the conditional convex order satisfies Axiom 8 on distributional invariance, $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ implies $(W,\mathbb{X})\preccurlyeq_{ccx}(W^{\prime},\mathbb{X}^{\prime})$. From Theorem 1.6, we then obtain $(F_{W}\circ F_{W|\mathbb{X}}^{-1}(V),V)\geq_{c}(F_{W^{\prime}}\circ F_{W^{\prime}|\mathbb{X}^{\prime}}^{-1}(V),V)$. Using that the concordance order is invariant under increasing transformations of the components and using the invariance in Proposition A.1ii, we obtain from Lemma B.6 that
$(F_{W|\mathbb{X}}^{-1}(V),F_{W}^{-1}(V))\geq_{sm}(F_{W^{\prime}|\mathbb{X}^{\prime}}^{-1}(V),F_{W^{\prime}}^{-1}(V)).$ (B.30)
Since the function $(y,y^{\prime})\mapsto-h(y^{\prime}-y)$ is supermodular whenever $h$ is convex, (B.30) yields
$\begin{split}\int_{\mathbb{R}^{p}}\mathcal{W}_{c}(\pi_{\mathbb{x}},\nu)\mathsf{\,d}\mu(\mathbb{x})&=\mathbb{E}h(F_{W|\mathbb{X}}^{-1}(V)-F_{W}^{-1}(V))\\ &\leq\mathbb{E}h(F_{W^{\prime}|\mathbb{X}^{\prime}}^{-1}(V)-F_{W^{\prime}}^{-1}(V))=\int_{\mathbb{R}^{p^{\prime}}}W_{c}(\pi^{\prime}_{\mathbb{x}},\nu^{\prime})\mathsf{\,d}\mu^{\prime}(\mathbb{x}),\end{split}$
where we use (32) for the first and, similarly, for the second equality. Here, $\nu^{\prime}$ and $\mu^{\prime}$ denote the distribution of $W^{\prime}$ and $\mathbb{X}^{\prime}$, respectively, and $\pi^{\prime}_{\mathbb{x}}$ refers to the conditional distribution $W^{\prime}\mid\mathbb{X}^{\prime}=\mathbb{x}$.
To prove statement i, let $Y$ and $\mathbb{X}$ be independent. Then $\pi_{\mathbb{x}}=\nu$ for $\mu$-almost all $\mathbb{x}\in\mathbb{R}^{p}$. Hence, $\mathcal{W}_{c}(\pi_{\mathbb{x}},\nu)=\mathcal{W}_{c}(\nu,\nu)=0$ for $\mu$-almost all $\mathbb{x}$ and thus $\overrightarrow{\mathcal{W}}(Y,\mathbb{X})=0$. Since $\overrightarrow{\mathcal{W}}$ is $\preccurlyeq_{ccx}$-increasing due to (33) and since independent random vectors are minimal elements in the conditional convex order due to Theorem 1.4, we obtain $\overrightarrow{\mathcal{W}}(Y^{\prime},\mathbb{X}^{\prime})\geq 0$ for all $(Y^{\prime},\mathbb{X}^{\prime})$. It remains to show that $\overrightarrow{\mathcal{W}}(Y^{\prime},\mathbb{X}^{\prime})\leq 1$. But this is a consequence of
$\int_{\mathbb{R}^{p}}\mathcal{W}_{c}(\pi_{\mathbb{x}},\nu)\mathsf{\,d}\mu(\mathbb{x})$ $\leq\int_{\mathbb{R}^{p}}\int_{\mathbb{R}}\int_{\mathbb{R}}c(y,y^{\prime})\mathsf{\,d}\pi_{\mathbb{x}}(y)\mathsf{\,d}\nu(y^{\prime})\mathsf{\,d}\mu(\mathbb{x})$ (B.31)
$=\int_{\mathbb{R}}\int_{\mathbb{R}}c(y,y^{\prime})\mathsf{\,d}\nu(y)\mathsf{\,d}\nu(y^{\prime}).$
To prove in statement ii the characterization of independence, it remains to show that $\overrightarrow{\mathcal{W}}(Y,\mathbb{X})=0$ implies independence of $Y$ and $\mathbb{X}$. Therefore, let $W=F_{Y}(Y)$ and $W^{\prime}=F_{Y^{\prime}}(Y^{\prime})$ and consider
$0=\int_{\mathbb{R}^{p}}\mathcal{W}_{c}(\pi_{\mathbb{x}},\nu)\mathsf{\,d}\mu(\mathbb{x})=\mathbb{E}h(F_{W|\mathbb{X}}^{-1}(V)-F_{W}^{-1}(V))$ (B.32)
where the second equality follows with Lemma B.7. We show that $\zeta:=F_{W|\mathbb{X}}^{-1}(V)-F_{W}^{-1}(V)=0$ almost surely. Then $W$ and $\mathbb{X}$, and thus $Y$ and $\mathbb{X}$, are independent. To this end, assume that $P(\zeta\neq 0)>0$. Since $F_{W|\mathbb{X}}^{-1}(V)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}F_{W}^{-1}(V)$ due to Lemma B.4, we have $\mathbb{E}\zeta=0$. Using that $h$ is strictly convex in 0, it follows that $\mathbb{E}h(\zeta)>h(0)=0$. But this contradicts (B.32). The proof that $\overrightarrow{\mathcal{W}}_{c}(Y,\mathbb{X})=1$ characterizes perfect dependence of $Y$ on $\mathbb{X}$ follows in the same line as the proof of [68, Theorem 2.2(iii)] noting that $h(y^{\prime}-y)+h(y-y^{\prime})>h(y-y)+h(y^{\prime}-y^{\prime})$ for $y\neq y^{\prime}$ since $h$ is convex and strictly convex in 0.
Statement iii on information monotonicity is a direct consequence of (33), together with information monotonicity of the conditional convex order.
To prove statement iv, consider the differences $\zeta:=F_{F_{Y}(Y)|\mathbb{X}}^{-1}(V)-F_{F_{Y}(Y)}^{-1}(V)$ and $\zeta^{\prime}:=F_{F_{Y}(Y)|(\mathbb{X},\mathbb{Z})}^{-1}(V)-F_{F_{Y}(Y)}^{-1}(V)$ for $V\sim U(0,1)$ independent of $(\mathbb{X},\mathbb{Z})$. Then, by Lemma B.8, we have $\zeta\leq_{cx}\zeta^{\prime}$. Hence, by Strassen’s theorem, there exist versions $\tilde{\zeta}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}\zeta$ and $\tilde{\zeta}^{\prime}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}\zeta^{\prime}$ such that $\mathbb{E}[\tilde{\zeta}^{\prime}\mid\tilde{\zeta}]=\tilde{\zeta}$ almost surely; see [51, Corollary 3.36]. Since $h$ is convex, we obtain for $\alpha:=\int_{\mathbb{R}}\int_{\mathbb{R}}c(y,y^{\prime})\mathsf{\,d}\nu(y)\mathsf{\,d}\nu(y^{\prime})$ that
[…]
$\leq\int_{\mathbb{R}}\mathbb{E}[h(\tilde{\zeta}^{\prime})\mid\tilde{\zeta}=w]\mathsf{\,d}P^{\tilde{\zeta}}(w)=\mathbb{E}h(\tilde{\zeta}^{\prime})=\alpha\,\overrightarrow{\mathcal{W}}_{c}(Y,(\mathbb{X},\mathbb{Z})),$ (B.33)
where the first and last equality are due to (32) using that $(y,y^{\prime})\mapsto h(y^{\prime}-y)$ is submodular because $h$ is convex. The above inequality is due to Jensen’s inequality for conditional expectations. Now, assume that $Y$ and $\mathbb{Z}$ are conditionally independent given $\mathbb{X}$. Then we have $\zeta\stackrel{{\scriptstyle\mathrm{d}}}{{=}}\zeta^{\prime}$. Hence, using the martingale property, the conditional distribution $\tilde{\zeta}^{\prime}\mid\tilde{\zeta}=w$ is degenerate for $P^{\zeta}$-almost all $w$ and the inequality in (B.33) becomes an equality. Conversely, if $Y$ and $\mathbb{Z}$ are not conditionally independent given $\mathbb{X}$, then there exists a Borel set $A$ with $P(\tilde{\zeta}\in A)>0$ such that $\tilde{\zeta}^{\prime}\mid\tilde{\zeta}=w$ is non-degenerate for all $w\in A$. Then, for $h$ strictly convex, the inequality in (B.33) is strict. $\blacksquare$
@@
(35): Due to Theorem 1.6, we know that $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ implies $(F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V),U)\leq_{c}(F_{Y^{\prime}}\circ q_{Y^{\prime};\mathbb{X}^{\prime}}^{\uparrow U}(V),U)$ and thus $\mathsf{R}_{\mu}(Y,\mathbb{X})=\mu(F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V),U)\leq\mu(F_{Y^{\prime}}\circ q_{Y^{\prime};\mathbb{X}^{\prime}}^{\uparrow U}(V),U)=\mathsf{R}_{\mu}(Y^{\prime},\mathbb{X}^{\prime})$.
@@
i & ii: $\mathbb{X}$ and $Y$ are independent if and only if $q_{Y;\mathbb{X}}^{\uparrow U}(V)$ and $U$ are independent; see Remark 1.7iii. Since $\mu$ characterizes independence on $\mathcal{R}^{\uparrow}$, the latter is equivalent to $\mathsf{R}_{\mu}(Y,\mathbb{X})=\mu(F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V),U)=0$. $Y$ perfectly depends on $\mathbb{X}$ if and only if $q_{Y;\mathbb{X}}^{\uparrow U}(V)$ and $U$ are comonotone; see Remark 1.7iii. Since $\mu$ characterizes comonotonicity on $\mathcal{R}^{\uparrow}$, the latter is equivalent to $\mathsf{R}_{\mu}(Y,\mathbb{X})=\mu(F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V),U)=1$. Assertion i then follows from (35).
iii: By Theorem 1.4, we have $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y,(\mathbb{X},\mathbb{Z}))$. Hence, (35) implies the statement.
@@
iv: Assume that $\mathsf{R}_{\mu}(Y,\mathbb{X})=\mathsf{R}_{\mu}(Y,(\mathbb{X},\mathbb{Z}))$. By information monotonicity of $\preccurlyeq_{ccx}$ and due to Theorem 1.6, we know that $(F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V),U)\leq_{c}(F_{Y}\circ q_{Y;(\mathbb{X},\mathbb{Z})}^{\uparrow U}(V),U)$. This implies $\mathsf{R}_{\mu}(Y,\mathbb{X})=\mu(F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V),U)\leq\mu(F_{Y}\circ q_{Y;(\mathbb{X},\mathbb{Z})}^{\uparrow U}(V),U)=\mathsf{R}_{\mu}(Y,(\mathbb{X},\mathbb{Z}))=\mathsf{R}_{\mu}(Y,\mathbb{X})$ and thus $\mu(F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V),U)=\mu(F_{Y}\circ q_{Y;(\mathbb{X},\mathbb{Z})}^{\uparrow U}(V),U)$. Since $\mu$ is strictly increasing on $\mathcal{R}^{\uparrow}$, it follows that $(F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V),U)=_{c}(F_{Y}\circ q_{Y;(\mathbb{X},\mathbb{Z})}^{\uparrow U}(V),U)$ and thus, by Theorem 1.6, $(Y,\mathbb{X})=_{ccx}(Y,(\mathbb{X},\mathbb{Z}))$. However, the latter means that $Y$ and $\mathbb{Z}$ are conditionally independent given $\mathbb{X}$. For the reverse direction, assume that $Y$ and $\mathbb{Z}$ are conditionally independent given $\mathbb{X}$. Then $(Y,\mathbb{X})=_{ccx}(Y,(\mathbb{X},\mathbb{Z}))$ and thus, by Theorem 1.6, $(F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V),U)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(F_{Y}\circ q_{Y;(\mathbb{X},\mathbb{Z})}^{\uparrow U}(V),U)$. This implies $\mathsf{R}_{\mu}(Y,\mathbb{X})=\mu(F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V),U)=\mu(F_{Y}\circ q_{Y;(\mathbb{X},\mathbb{Z})}^{\uparrow U}(V),U)=\mathsf{R}_{\mu}(Y,(\mathbb{X},\mathbb{Z}))$. $\blacksquare$
@@ B.7 Proofs of Section 5
$(q_{Y;X}^{\uparrow U}(V),U)\leq_{c}(q_{Y^{\prime};X^{\prime}}^{\uparrow U}(V),U)$ (B.34)
@@
$q_{Y;X}^{\uparrow U}(V)=F_{U}^{-1}(V)=F_{Y|X=q_{X}(U)}^{-1}(V),$ (B.35)
with $q_{Y;X}^{\uparrow u}(v)$ defined by (13). Similarly, we obtain $q_{Y^{\prime};X^{\prime}}^{\uparrow U}(V)=F_{Y^{\prime}|X^{\prime}=q_{X^{\prime}}(U)}^{-1}(V)$. Now, observe that
$(Y,X)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(F_{Y|X=q_{X}(U)}^{-1}(V),q_{X}(U))\quad\text{and}\quad(Y^{\prime},X^{\prime})\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(F_{Y^{\prime}|X^{\prime}=q_{X^{\prime}}(U)}^{-1}(V),q_{X^{\prime}}(U));$ (B.36)
see (A.3). Since the concordance order is invariant under increasing transformations, we obtain from $q_{X}=q_{X^{\prime}}$ with (B.34) and (B.35) that $(Y,X)\leq_{c}(Y^{\prime},X^{\prime})$.
The reverse direction is a special case of Theorem 3.3ii. $\blacksquare$